How many liters of a 70% alcohol solution must be mixed with 30 liters of a 15% alcohol solution to obtain a solution that is 60%?
Answer by Earlsdon(6294)

Let x = the required number of liters of 70% alcohol solution, change the percentages to their decimal equivalents, then...
0.7x+(0.15)(30) = (30+x)(0.6)
0.7x+4.5 = 18+0.6x Subtract 0.6x from both sides.
0.1x+4.5 = 18 Subtract 4.5 from both sides.
0.1x = 13.5 Finally, divide both sides by 0.1
x = 135 liters.
You will need to mix 135 liters of 70% alcohol solution with 30 liters of 15% alcohol solution to obtain 165 liters of 60% alcohol solution.